If the light ray moves from inside the glass toward the glass-air interface at an angle of 35.0° to the normal, determine the angle of refraction. The ray bends ________° away from the normal, as expected.


1
Expert's answer
2020-07-07T10:02:44-0400

The angle of refraction "\theta_2" is given by the Snell's law:


"\sin \theta_2 = \dfrac{n_1}{n_2}\sin\theta_1"

where "\theta_1 = 35\degree" is the angle of incidence, "n_1 = 1.52" is the refractive ingex of glaass, "n_2 = 1" is the refractive ingex of air. Thus


"\sin \theta_2 = \dfrac{1.52}{1}\sin35\degree \approx 0.87\\
\theta_2 \approx \arcsin0.87 \approx 60.5\degree"

Answer. The ray bends 60.5° away from the normal, as expected.